# Graph Neural Networks for Virtual Sensing in Complex Systems: Addressing Heterogeneous Temporal Dynamics

### v1

@@ 6.2.2 In-Depth Performance Evaluation
Impact of temperature on signal characteristics: To further investigate the influence of temperature on bridge dynamics and its implications for load prediction, we analyze the frequency characteristics of acceleration signals under varying temperatures. Fig. 11(a) shows raw acceleration signals at -7.3°C and 29.2°C, revealing notable differences in signal amplitude and pattern. This temperature-dependent shift in frequency content is further analyzed in Fig. 11(b), where the spectral centroid (a measure of the dominant frequency) exhibits a positive correlation with temperature. This finding suggests that higher temperatures lead to an increase in the dominant frequency of acceleration signals. Such variations in signal characteristics could contribute to the observed decrease in prediction accuracy for all models at lower temperatures, likely due to the model’s inherent limitations in capturing the complex non-linear relationship between temperature and the frequency response of the bridge structure.
Effectiveness of the Gating Mechanism in 1D-GCNN. During the colder months, 1D-GCNN performs noticeably better than 1D-CNN, highlighting the effectiveness of the gating mechanism using temperature as a gating signal. This approach helps to select frequency components more relevant to the temperature conditions, demonstrating the benefits of incorporating contextual information into the model architecture. This contextual adaptation allows 1D-GCNN to maintain higher prediction accuracy by dynamically adjusting to the temperature-dependent variations in the acceleration signals.
@@ 7 Conclusion
In this research, we have presented HTGNN, a novel GNN-based virtual sensor framework specifically designed to address the challenges of heterogeneous temporal dynamics and varying operating conditions in complex systems. Through extensive experiments on both a test-rig bearing dataset, characterized by high-frequency vibrations and temperature fluctuations under varying rotational speeds, and a simulated train-bridge-track interaction dataset, capturing multi-scale displacement and acceleration data influenced by ambient temperature variations, we have demonstrated HTGNN’s superior performance and adaptability in virtual load sensing across diverse industrial domains, highlighting its potential for broad applicability. HTGNN’s effectiveness stems from its ability to explicitly model the complex, heterogeneous relationships between sensor modalities and its capacity to extract operating condition-aware dynamics. This approach allows HTGNN to adapt to changing operating conditions and accurately predict loads even in scenarios where traditional methods struggle, such as low rotational speeds for bearings or extreme temperatures for bridges. The consistent robustness and accuracy of HTGNN in both case studies highlight its potential as a reliable virtual sensor for diverse IIoT applications, enabling effective monitoring, predictive maintenance, and enhanced system performance. Our findings highlight the importance of leveraging physical prior and domain knowledge in the design of virtual sensors, particularly through the use of GNNs. By encoding domain knowledge and physical priors about system behavior into the construction of the graph structure, GNNs can explicitly model the physical relationships between sensors, providing the model with valuable insights into load distribution across the system and leading to significantly improved prediction accuracy. While this work focuses on virtual sensing for load prediction, the underlying principles of HTGNN, such as heterogeneous interaction modeling and operating condition-aware dynamics extraction, are generalizable and can be applied to other downstream tasks within the IIoT domain. For example, HTGNN could be adapted for tasks like signal reconstruction, anomaly detection, or forecasting, where the ability to handle heterogeneous temporal dynamics and varying operating conditions is crucial.
@@ 8 Data and Code Availability
The script we used to generate this synthetic dataset is available in the associated code repository to ensure reproducibility and facilitate further studies. Our code and data are available under https://github.com/EPFL-IMOS/htgnn.
@@ A.3 Hyperparameter Tuning
Table 4: Range of hyperparameters with the optimal values indicated within parentheses.
@@
Initial hyperparameters were drawn from relevant literature and subsequently refined through a grid search over an expanded parameter space tailored to the specific sensitivities of our case studies. Table 4 summarizes the optimal configurations for each case study. Values in parentheses indicate the best configuration for the respective case study; a single value denotes the optimal choice for both. An asterisk () indicates additional details provided in the text. The grid search explored variations in the number of layers (L), hidden layer dimensions (H), normalization techniques (Batch Normalization (BN) and Layer Normalization (LN)), and dropout rates $p$ (ranging from 0 to 0.5). The hyperparameter search space was designed to maintain a roughly consistent number of parameters across all models. Model-specific parameters and optimal model sizes are also detailed in Table 4. The final hyperparameter selections were based on achieving the lowest validation loss. Specifically, the hyperparameter tuning process for each baseline model is as follows:

### v2

@@ 6.2.2 In-Depth Performance Evaluation
Impact of temperature on signal characteristics: To further investigate the influence of the exogenous variable, temperature, on bridge dynamics and its implications for load prediction, we analyze the frequency characteristics of acceleration signals under varying temperatures. Fig. 11(a) shows raw acceleration signals at -7.3°C and 29.2°C, revealing notable differences in signal amplitude and pattern. This temperature-dependent shift in frequency content is further analyzed in Fig. 11(b), where the spectral centroid (a measure of the dominant frequency) exhibits a positive correlation with temperature. This finding suggests that higher temperatures lead to an increase in the dominant frequency of acceleration signals. Such variations in signal characteristics could contribute to the observed decrease in prediction accuracy for all models at lower temperatures, likely due to the model’s inherent limitations in capturing the complex non-linear relationship between temperature and the frequency response of the bridge structure.
Effectiveness of the Gating Mechanism in 1D-GCNN. During colder months, 1D-GCNN performs noticeably better than 1D-CNN, highlighting the effectiveness of the gating mechanism using temperature as a gating signal. This approach helps select frequency components more relevant to temperature conditions, demonstrating the benefits of incorporating contextual information into the model architecture. This contextual adaptation allows 1D-GCNN to maintain higher prediction accuracy by dynamically adjusting to the temperature-dependent variations in the acceleration signals.
Limitation of simple multiscale modeling. While TimeMixer’s multiscale mixing strategy effectively captures temporal dependencies at different resolutions, it fails to explicitly model the heterogeneous interactions between these scales, particularly under strong exogenous influences. This limitation is particularly evident in the bridge dataset, where temperature significantly affects both acceleration and displacement signals. TimeMixer’s inability to capture these interactions and the exogenous influence on different scales likely contributes to its deteriorated performance at lower temperatures.
@@ 7 Conclusion
In this research, we have presented HTGNN, a novel GNN-based virtual sensor framework specifically designed to address the challenges of heterogeneous temporal dynamics and varying operating conditions in complex systems. Through extensive experiments on both a test-rig bearing dataset, characterized by high-frequency vibrations and temperature fluctuations under varying rotational speeds, and a simulated train-bridge-track interaction dataset, capturing multi-scale displacement and acceleration data influenced by ambient temperature variations, we have demonstrated HTGNN’s superior performance and adaptability in virtual load sensing across diverse industrial domains, highlighting its potential for broad applicability. HTGNN’s effectiveness stems from its ability to explicitly model the complex, heterogeneous relationships between sensor modalities and its capacity to extract operating condition-aware dynamics. This approach allows HTGNN to adapt to changing operating conditions and accurately predict loads even in scenarios where traditional methods struggle, such as low rotational speeds for bearings or extreme temperatures for bridges. The consistent robustness and accuracy of HTGNN in both case studies highlight its potential as a reliable virtual sensor for diverse IIoT applications, enabling effective monitoring, predictive maintenance, and enhanced system performance. Our findings highlight the importance of leveraging physical prior and domain knowledge in the design of virtual sensors, particularly through the use of GNNs. By encoding domain knowledge and physical priors about system behavior into the construction of the graph structure, GNNs can explicitly model the physical relationships between sensors, providing the model with valuable insights into load distribution across the system and leading to significantly improved prediction accuracy. Furthermore, HTGNN demonstrates promising computational efficiency and scalability. By sharing encoders for all nodes of the same type, the model can compute node embeddings in parallel. This parallel processing, combined with the use of efficient GNN operations, significantly reduces the overall computation time. Consequently, HTGNN can efficiently handle large-scale industrial systems with many sensors. The model’s ability to perform fast inference enables real-time load prediction, which is crucial for applications such as real-time monitoring and control in industrial settings. While this work focuses on virtual sensing for load prediction, the underlying principles of HTGNN, such as heterogeneous interaction modeling and operating condition-aware dynamics extraction, are generalizable and can be applied to other downstream tasks within the IIoT domain. For example, HTGNN could be adapted for tasks like signal reconstruction, anomaly detection, or forecasting, where the ability to handle heterogeneous temporal dynamics and varying operating conditions is crucial.
@@ 8 Data and Code Availability
The bearing test rig dataset is available at https://doi.org/10.5281/zenodo.14959001, and the bridge dataset can be accessed at https://doi.org/10.5281/zenodo.14972955. The code repository at https://github.com/EPFL-IMOS/htgnn contains data loaders for both datasets, along with the implementation of the proposed method and the baseline methods.
@@ A.3 Hyperparameter Tuning
Table 4: Range of hyperparameters with optimal values indicated within parentheses.
@@
Initial hyperparameters, including the selection of data preprocessing parameters (window size and stride), were drawn from relevant literature and subsequently refined through a grid search over an expanded parameter space tailored to the specific sensitivities of our case studies. Data preprocessing parameters: For each dataset, the window size and stride used for data preprocessing were empirically determined. To ensure a fair comparison across all methods, these values were determined based on preliminary experiments with the baseline 1D-CNN model and then kept consistent for all other baseline models as well as our proposed method. For the window size, we tested values of 30 and 60. The optimal size varied between case studies, likely due to differences in sampling frequency and the nature of the underlying dynamic. Specifically, a window size of 60 was optimal for the bearing case study, while 30 was optimal for the train-track-bridge case study. The stride was also adjusted per dataset. We used a stride of 1 for the bearing case study and 5 for the train-track-bridge case study, reflecting the larger size of the latter. Model hyperparameters: Table 4 summarizes the optimal model configurations for each case study. Values in parentheses indicate the best configuration for the respective case study; a single value denotes the optimal choice for both. An asterisk () indicates additional details provided in the text. The grid search explored variations in the number of layers (L), hidden layer dimensions (H), normalization techniques (Batch Normalization (BN) and Layer Normalization (LN)), and dropout rates $p$ (ranging from 0 to 0.5). The hyperparameter search space was designed to maintain a roughly consistent number of parameters across all models. Model-specific parameters and optimal model sizes are also detailed in Table 4. The final hyperparameter selections were based on achieving the lowest validation loss. Specifically, the hyperparameter tuning process for each baseline model is as follows:
@@
  • TimeMixer: Adapted from [35], we varied encoder layers (2-6). The hidden dimension was varied from 16-64, with the final fully connected layer’s dimension set as double. Averaging was the best downsampling strategy for the bearing case study and convolution for the bridge case study. For both, 2 downsampling layers (1-3 tested) were optimal. Moving average decomposition was best for the bearing dataset, while Discrete Fourier Transform (DFT) decomposition with 9 top frequencies (3-15 tested) was best for the bridge dataset. Reversible Instance Normalization (RevIN) [67] normalization was not found to be useful.